OpenStudy (anonymous):

what is the lcd that is going to eliminate the fractions in the equation , 4y/5-7/4+7/10=7/2?

OpenStudy (anonymous):

\(\frac{7}{4}+\frac{7}{10}=\frac{7}{2}\) Start by finding the prime factors of each denominator.